# Explainable Anomaly Detection in Images and Videos: A Survey

### v2

@@ VIII Conclusion
This paper presents the first survey on explainable anomaly detection for 2D visual data. The explainable and interpretable anomaly detection approaches for both images and videos are thoroughly summarized and how these methods explain the anomaly identification process is well elaborated. The reasons why certain explainable anomaly detection methods are applicable to both images and videos, while others are limited to a single modality are further explored. Moreover, summaries of current datasets and evaluation metrics for 2D visual anomaly detection are offered. The survey is closed by pointing out some future directions for a deeper and better understanding of anomaly detection in images and videos.

### v4

@@ VIII Conclusion
This paper presents the first survey focused specifically on explainable anomaly detection for 2D visual data (images and videos). We provided a comprehensive overview of the growing body of work in explainable image anomaly detection (IAD) and explainable video anomaly detection (VAD), categorizing methods based on their core techniques, including attention, input perturbation, generative models, reasoning, intrinsic attributes, and foundation models. We elaborated on how these methods aim to explain the anomaly identification process and analyzed the factors influencing their applicability across image and video modalities. Furthermore, we summarized relevant datasets, highlighting their suitability for different types of anomalies and explanation evaluations, and discussed evaluation metrics, covering both standard performance measures and crucial aspects of explanation quality like faithfulness and utility. The survey concludes by identifying key open challenges and promising future directions.